# Validation tests of Gaussian boson samplers with photon-number resolving detectors

### v1

@@ III Normally-ordered phase-space methods
Here we summarize phase-space methods, and explain the motivation for using them. The purpose is to transform an exponentially complex Hilbert space into a positive distribution that can be randomly sampled. Normally-ordering methods are the closest to experimental statistics using photon-counting. Representations that are not normally ordered, like the Wigner [37] or Q-function [38] approach, have a rapidly growing sampling error with system size.
@@ III.1 Glauber-Sudarshan representation
hence its commonly called the diagonal P-representation, where $\left|\boldsymbol{\alpha}\right\rangle=\bigotimes_{j=1}^{M}\left|\alpha_{j}\right\rangle$ is a multi-mode coherent state vector with eigenvalues $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{M}]$ and $P(\boldsymbol{\alpha})$ is the P-representation probability distribution. The diagonal P-representation can be used to evaluate expectation values of the product of normally ordered operators due to the relationship
@@
$\mathcal{G}_{S}^{(M)}(m)=\int P(\boldsymbol{\alpha})\left[\frac{1}{m!}\left(n^{\prime}_{S}\right)^{m}e^{-n^{\prime}_{S}}\right]\text{d}^{2M}\boldsymbol{\alpha},$ (14)
where $S=\{1,\dots,M\}$. However, this is only valid for classical states such as coherent, thermal and squashed states [39] as they are defined entirely by their diagonal density matrix elements. Since the density operator of nonclassical states contains off-diagonal elements due to quantum superpositions, if the diagonal P-representation is used for simulations of nonclassical states $P(\boldsymbol{\alpha})$ becomes singular and negative. From standard mathematical requirements that it should be real, normalizable, non-negative and non-singular, it is no longer a probability distribution.
@@ III.2 Positive-P representation
where the additional dimensions allow off-diagonal density matrix elements to be included in the coherent state basis $\left|\boldsymbol{\beta}\right\rangle=\bigotimes_{j=1}^{M}\left|\beta_{j}\right\rangle$ with eigenvalues $\boldsymbol{\beta}=[\beta_{1},\dots,\beta_{M}]$. The coherent amplitudes $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$ are independent and can vary along the entire complex plane. We can also take the real part of Eq.(15), since the density matrix is hermitian. The benefit of using the positive P-representation for photon counting experiments is that it includes the diagonal P-representation when the positive-P distribution satisfies $P(\boldsymbol{\alpha},\boldsymbol{\beta})=P(\boldsymbol{\alpha})\delta(\text{$\boldsymbol{\alpha}$-$\boldsymbol{\beta}$})$. This allows us to define classical phase-space as having $\boldsymbol{\beta}=\boldsymbol{\alpha}$, while a nonclassical phase space requires non-vanishing probability of $\boldsymbol{\beta}\neq\boldsymbol{\alpha}$.
@@
We have taken the real part as is required for the applications presented in the next section.
@@
Unlike methods for threshold detectors [15], Eq.(18) doesn’t require multidimensional inverse discrete Fourier transforms, which increase computation time significantly [16]. Since threshold detector outputs are always binary, the largest number of observable photon counts in a pattern corresponds to a detection event in every output mode, such that $\sum_{i}^{M}c_{i}\leq M$ where the largest observable grouped count is $m^{(M)}$.
Since PNR detectors can resolve multiple photons arriving at each detector, an experiment or classical algorithm replicating photo-detection measurements can now produce count patterns where $\sum_{i}^{M}c_{i}>M$, in which case $m^{(M)}$ is no longer the largest observable grouped count. Instead of scaling with mode number, numerical simulations scale with photon count bins, where we define the maximum observable grouped count as $m^{(\text{max})}$. For experiments with small mean input photon numbers or high losses, one may have $m^{(\text{max})}<m^{(M)}$, while low losses and large photon numbers will produce $m^{(\text{max})}>m^{(M)}$. Typically, we choose $m^{(\text{max})}$ to occur when probabilities are of the order $\mathcal{G}_{S}^{(n)}(m^{(\text{max})})=10^{-7}$ or smaller, providing a standard cutoff for very small probabilities as was implemented in previous methods [15, 16].
Since our observable of interest is the total output photon number Eq.(9), this method requires a photon counting experiment to accurately measure the number of counts arriving at each detector, such that $\left\langle\hat{p}_{j}(c_{j}^{(\text{max})}+1)\right\rangle=0$. Failure to satisfy this requirement would mean output photons in the summations Eq.(9) would be larger than detected counts, leading to inaccurate comparisons. In some regards this is a benefit, as it allows one to determine whether the experimental counts are actually measuring all possible output photons, as some exact algorithms exploit these unmeasured photons to decrease GBS simulation time [20].
@@ III.3 Squeezed states in phase-space
For small $\epsilon$, thermalized squeezed states remain nonclassical and produce a modified output distribution that more closely resembles recent experimental distributions using threshold detectors [15, 16]. Due to this non-classicality, it has been shown [16] that even if an experiments claim of quantum advantage against the ideal distribution becomes invalidated by classical algorithms that replicate photo-detection measurements [18, 19, 16], it may still be able to claim quantum advantage against this non-ideal thermalized distribution.
From this model of decoherence, one can introduce a general random sampling formalism to generate initial stochastic samples for any Gaussian input [15]
@@
where $\left\langle w_{j}w_{k}\right\rangle_{P}=\delta_{jk}$ are real Gaussian noises and the quadrature variances are defined in Eq.(3). For classical inputs, $\sigma_{y_{j}}$ is real and $\alpha_{j}=\beta_{j}$. For quantum inputs, $\sigma_{y_{j}}$ is imaginary and $\alpha_{j}\neq\beta_{j}$. Following from Eq.(4), the output stochastic amplitudes used to perform simulations are then obtained via the transformations $\boldsymbol{\alpha}^{\prime}=\boldsymbol{T}\boldsymbol{\alpha}$ and $\boldsymbol{\beta}^{\prime}=\boldsymbol{T}\boldsymbol{\beta}$ where the amplitudes $\boldsymbol{\alpha}^{\prime}$, $\boldsymbol{\beta}^{\prime}$ correspond to measurements of the density matrix $\hat{\rho}^{\text{(out)}}$.
@@ IV Exact models of GBS grouped count distributions
In this section, we focus on comparing positive-P moments to exact distributions of GBS networks whose size, mean squeezing parameters and loss rates correspond to those found in the PNR experiments of Madsen et al [5]. However, we note that when even-odd count oscillations are present in lossless and ultra-low loss photonic networks, the positive-P moments suffer from slow convergence and large sampling errors.
Since the even-odd count oscillations vanish as soon as one has a high chance of losing a photon, the lossless and ultra-low loss regimes are not relevant to any current or near future experimental network, which are both far from lossless and whose photon number distribution is approximately Gaussian for large system sizes with loss rates as small as $t=0.95$. Therefore, a detailed theoretical analysis of the positive P-representation in the lossless and ultra-low loss regimes will not be performed here, as projected methods [21] may converge faster in such regimes.
IV.1 Comparisons with small and large sized networks
@@
Comparisons between positive-P moments and the exact distribution Eq.(24) are presented in Fig.(1a) and Fig.(1b) for these networks where simulations are performed for an ensemble size of $E_{S}=2.4\times 10^{6}$. In the limit of large $E_{S}$ the positive-P moments converge to moments of a distribution (see Eq.(16)). The exact size of $E_{S}$ required for a given accuracy varies. An ensemble size of the order $E_{S}\geq 10^{6}$ is generally needed for relative errors less than $10^{-3}$, typical of current experimental accuracy.
@@
Figure 1: Exact multi-mode squeezed state photon counting distributions (dashed black lines) for various uniform amplitude loss rates $t$ are compared to positive-P moments of the GCP Eq.(10) (solid blue lines) for various network sizes, squeezing parameters and loss rates corresponding to the various experimental networks implemented in Ref.[5]. a) A GBS set-up with $M=N=16$ is considered where pure squeezed states with uniform squeezing parameter $\boldsymbol{r}=[0.89,\dots,0.89]$ are transformed by a Haar random unitary matrix multiplied by a uniform amplitude loss coefficient $t\boldsymbol{U}$ with $t=0.6$. Network sizes are now increased to b) $M=N=72$, with $\boldsymbol{r}=[0.89,\dots,0.89]$ and $t=0.56$, c) $M=N=216$ with $\boldsymbol{r}=[1.1,\dots,1.1]$ and $t=0.57$, and d) $M=N=288$ mode network with $\boldsymbol{r}=[1,\dots,1]$ and $t=0.6$. In all comparisons, the even-odd grouped count oscillations characteristic of pure squeezed photon statistics are not present as losses are too large. Positive-P moments are obtained by averaging over sample ensembles of size $E_{S}=2.4\times 10^{6}$ where upper and lower lines correspond to ±$\pm 1\sigma_{T,j}$, where $\sigma_{T,j}\propto 1/\sqrt{E_{S}}$ is the theoretical sampling errors for the $j$-th photon count bin.
@@ V.1 Statistical tests
Here we use the shorthand notation $\mathcal{\bar{G}}_{i}=\left\langle\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right\rangle_{P}=E_{S}^{-1}\sum_{e=1}^{E_{S}}\left(\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right)_{(e)}$ for the stochastic trajectory $e\in E_{S}$ to denote the positive-P GCP ensemble means of the $i=1,2,\dots,k$ photon count bin for the $j=1,\dots,d$ grouped count, which coverages to the true theoretical GCP in the limit $\mathcal{G}_{S,i}=\lim_{E_{S}\rightarrow\infty}\bar{\mathcal{G}}_{S,i}$ for any general input state $S$. Simulated ensemble means are then compared to the experimental GCPs $\mathcal{G}_{E,i}=N_{E}^{-1}\sum_{i}c_{i}$ formed from $N_{E}\subset\mathcal{S}_{P}$ photon count patterns.
@@
Chi-square tests are used for two reasons: the first is that its a standard test applied to data from random number generators [14], where the summation is performed over $k$ count bins satisfying $\sum_{i\in S_{j}}c_{i}>10$. The second is that it contains an explicit dependence on both theoretical and experimental sampling errors as $\sigma_{i}^{2}=\sigma_{E,i}^{2}+\sigma_{T,i}^{2}$, where $\sigma_{E,i}\propto 1/\sqrt{N_{E}}$ is the estimated experimental sampling error. This provides a method of accurately determining whether observed differences are caused by large sampling error, as is the case for sparse data.
@@
Output $\chi^{2}$ values follow a chi-squared distribution that is positively skewed for small $k$. Performing a Wilson-Hilferty transformation [45], which transforms the chi-square result as $\chi^{2}\rightarrow\left(\chi^{2}/k\right)^{1/3}$, the chi-square distribution converges to a Gaussian distribution $\left(\chi^{2}/k\right)^{1/3}\rightarrow\mathcal{N}\left(\mu_{\mathcal{N}},\sigma_{\mathcal{N}}^{2}\right)$ with mean $\mu_{\mathcal{N}}=(9k-2)/9k$ and variance $\sigma_{\mathcal{N}}^{2}=2/(9k)$ for $k\geq 10$ due to the central limit theorem [45, 46]. This allows one to convert chi-square results into measurable probability distances using approximate Z-statistic, or Z-score, tests [47]
$Z=\frac{\left(\chi^{2}/k\right)^{1/3}-\mu_{\mathcal{N}}}{\sigma_{\mathcal{N}}},$ (28)
which measures how many standard deviations away a test statistic is from its expected normally distributed mean.
The results presented in subsequent subsections use the notation introduced in [16]. Comparisons between theory and experiment for simulations with pure squeezed states input into the experimental $M\times M$ $\boldsymbol{T}$-matrix are denoted by the subscript $EI$, which we call the ideal ground truth distribution. Our ideal ground truth corresponds to the ground truth distributions used in Ref. [5]. However they differ from the non-ideal, or modified, ground truth distribution that includes realistic experimental effects such as thermalization, and measurement corrections in the theoretical simulations, which are denoted by the subscript $ET$.
@@
In all cases, agreement within sampling errors produces test results of $\chi^{2}/k\approx 1$ and $\left|Z\right|\leq 1$ when accounting for stochastic fluctuations in the output statistics. For Z-statistic tests we use a regime of validity with the limit $Z\leq 6$ to define accurate comparisons. Above this limit test statistics are more than $6\sigma_{\mathcal{N}}$ from their predicted means, corresponding to very small observable probabilities.
V.2 Grouped count probability comparisons
Due to claims of quantum advantage, in this section we focus predominantly on validating count patterns from the $216$-mode high squeezing (HS) parameter and $288$-mode experimental data sets. Summary tables of chi-square and Z-statistic test results for all Borealis data sets can be found in Appendix A.

### v2

@@ III Normally-ordered phase-space methods
Here we summarize phase-space methods, and explain the motivation for using them. The purpose is to transform an exponentially complex Hilbert space into a positive distribution that can be randomly sampled. Normal-ordering methods are the closest to experimental statistics using photon-counting. Representations that are not normally ordered, like the Wigner [38] or Q-function [39] approach, have a rapidly growing sampling error with system size.
@@ III.1 Glauber-Sudarshan representation
hence its commonly called the diagonal P-representation, where $\left|\boldsymbol{\alpha}\right\rangle=\bigotimes_{j=1}^{M}\left|\alpha_{j}\right\rangle$ is a multi-mode coherent state vector with eigenvalues $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{M}]^{T}$ and $P(\boldsymbol{\alpha})$ is the P-representation probability distribution. The diagonal P-representation can be used to evaluate expectation values of the product of normally ordered operators due to the relationship
@@
$\mathcal{G}_{S}^{(M)}(m)=\int P(\boldsymbol{\alpha})\left[\frac{1}{m!}\left(n^{\prime}_{S}\right)^{m}e^{-n^{\prime}_{S}}\right]\text{d}^{2M}\boldsymbol{\alpha}.$ (14)
Since the density operator of nonclassical states contains off-diagonal elements due to quantum superpositions, if the diagonal P-representation is used for simulations of nonclassical states, $P(\boldsymbol{\alpha})$ becomes singular and negative. From standard mathematical requirements that it should be real, normalizable, non-negative and non-singular, it is no longer a probability distribution. This method is therefore only useful for classical states such as coherent, thermal and squashed states [40, 41], since these have a positive diagonal P-distribution.
@@ III.2 Positive-P representation
where the additional dimensions allow off-diagonal density matrix elements to be included in the coherent state basis $\left|\boldsymbol{\beta}\right\rangle=\bigotimes_{j=1}^{M}\left|\beta_{j}\right\rangle$ with eigenvalues $\boldsymbol{\beta}=[\beta_{1},\dots,\beta_{M}]^{T}$. The coherent amplitudes $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$ are independent and can vary along the entire complex plane. We can also take the real part of Eq.(15), since the density matrix is hermitian. The benefit of using the positive P-representation for photon counting experiments is that it includes the diagonal P-representation when the positive-P distribution satisfies $P(\boldsymbol{\alpha},\boldsymbol{\beta})=P(\boldsymbol{\alpha})\delta(\text{$\boldsymbol{\alpha}$-$\boldsymbol{\beta}$})$. This allows us to define classical phase-space as having $\boldsymbol{\beta}=\boldsymbol{\alpha}$, while a nonclassical phase space requires non-vanishing probability of $\boldsymbol{\beta}\neq\boldsymbol{\alpha}$.
@@
In Eq.(18), the required summation over output photon number is incorporated into the pattern projector, allowing one to make the substitution $m_{j}=\sum_{i\in S_{j}}c_{i}$, while we take the real part as required for pure squeezed state applications presented in the next subsection.
@@
Unlike methods for threshold detectors [9], Eq.(18) doesn’t require multidimensional inverse discrete Fourier transforms, which can increase the computation time. Since threshold detector outputs are always binary, the largest number of observable photon counts in a pattern corresponds to a detection event in every output mode, such that $\sum_{i}^{M}c_{i}\leq M$ where the largest observable grouped count is $m^{(M)}$.
Since PNR detectors can resolve multiple photons arriving at each detector, an experiment or classical algorithm replicating photo-detection measurements can now produce count patterns where $\sum_{i}^{M}c_{i}>M$, in which case $m^{(M)}$ is no longer the largest observable grouped count. Instead of scaling with mode number, numerical simulations scale with photon count bins, where we define the maximum observable grouped count as $m^{(\text{max})}$. For experiments with small mean input photon numbers or high losses, one may have $m^{(\text{max})}<m^{(M)}$, while low losses and large photon numbers will produce $m^{(\text{max})}>m^{(M)}$.
Typically, we choose $m^{(\text{max})}$ to occur when probabilities are of the order $\mathcal{G}_{S}^{(n)}(m^{(\text{max})})\leq 10^{-7}$, providing a standard cutoff for very small probabilities as was implemented in previous methods [9, 17]. A minimum observable grouped count $m^{(\text{min})}$ can also be introduced to form an analogous cut-off $\mathcal{G}_{S}^{(n)}(m^{(\text{min})})\leq 10^{-7}$. This allows our algorithm to only simulate grouped counts within the relevant range $m_{j}^{(\text{min})}\leq m_{j}\leq m_{j}^{(\text{max})}$, reducing memory use and computation times.
@@ III.3 Squeezed states in phase-space
Thermalized squeezed states, defined by the $\epsilon$ component introduced in Section II, are readily incorporated into the positive-P distribution. A general random sampling formalism is used to generate initial stochastic samples for any Gaussian input [9]
@@
where $\left\langle w_{j}w_{k}\right\rangle_{P}=\delta_{jk}$ are real Gaussian noises and the variances $\sigma_{x_{j}}$, $\sigma_{y_{j}}$ are defined in Eq.(3). For classical inputs, $\sigma_{y_{j}}$ is real and $\alpha_{j}=\beta_{j}$. For quantum inputs, $\sigma_{y_{j}}$ is imaginary and $\alpha_{j}\neq\beta_{j}$. Following from Eq.(4), the output stochastic amplitudes used to perform simulations are then obtained via the transformations $\boldsymbol{\alpha}^{\prime}=\boldsymbol{T}\boldsymbol{\alpha}$ and $\boldsymbol{\beta}^{\prime}=\boldsymbol{T}\boldsymbol{\beta}$ where the amplitudes $\boldsymbol{\alpha}^{\prime}$, $\boldsymbol{\beta}^{\prime}$ correspond to measurements of the density matrix $\hat{\rho}^{\text{(out)}}$.
@@ IV Exact models of GBS grouped count distributions
In this section, we validate the sampled positive-P method by comparing the results to exactly known distributions of GBS networks whose size, mean squeezing parameters and average loss rates correspond to those found in the PNR experiments of Madsen et al [5]. This demonstrates that for the relevant regime, the phase-space sampling errors are negligible. A full comparison with experiment that uses the inhomogeneous parameters describing the squeezing vector and transmission matrix is carried out in Section V.
The positive-P algorithm used throughout this paper is part of the phase-space network simulation program xqsim, which is available from the Github repository [45].
IV.1 Exactly known distributions
@@
When even-odd count oscillations are present in lossless and ultra-low loss photonic networks, the positive-P moments suffer from slow convergence and large sampling errors. Since the even-odd count oscillations vanish as soon as one has a high chance of losing a photon, the lossless and ultra-low loss regimes are not relevant to any current or near future experimental network, which are both far from lossless and whose photon number distribution is approximately Gaussian for large system sizes with loss rates as small as $t=0.95$. Therefore, a detailed theoretical analysis of the positive P-representation in the lossless and ultra-low loss regimes will not be performed here, as projected methods [22] may converge faster in such regimes.
IV.2 Comparisons with small and large sized networks
@@
Comparisons between positive-P moments and the exact distribution Eq.(24) are presented in Fig.(1a) and Fig.(1b) for these networks where simulations are performed for an ensemble size of $E_{S}=2.4\times 10^{6}$. In the limit of large $E_{S}$ the positive-P moments converge to moments of a distribution (see Eq.(16)). The exact size of $E_{S}$ required for a given accuracy varies. An ensemble size of the order $E_{S}\geq 10^{6}$ is generally needed for sampling and difference errors of less than $10^{-3}$, typical of current experimental accuracy.
@@
Figure 1: Validation of the positive-P GCP distributions (solid blue lines) by comparison to exact multi-mode squeezed state photon counting distributions (dashed black lines) in an exactly soluble example. We use uniform amplitude loss rates $t$ for different network sizes, squeezing parameters and loss rates corresponding to the various experimental networks implemented in Ref.[5]. a) A GBS set-up with $M=N=16$ is considered where pure squeezed states with uniform squeezing parameter $\boldsymbol{r}=[0.89,\dots,0.89]$ are transformed by a Haar random unitary matrix multiplied by a uniform amplitude loss coefficient $t\boldsymbol{U}$ with $t=0.6$. Network sizes are now increased to b) $M=N=72$, with $\boldsymbol{r}=[0.89,\dots,0.89]$ and $t=0.56$, c) $M=N=216$ with $\boldsymbol{r}=[1.1,\dots,1.1]$ and $t=0.57$, and d) $M=N=288$ with $\boldsymbol{r}=[1,\dots,1]$ and $t=0.6$. Positive-P distributions are obtained by averaging over the ensemble of samples $E_{S}=2.4\times 10^{6}$. Upper and lower lines correspond to ±$\pm 1\sigma_{T,j}$, where $\sigma_{T,j}\propto 1/\sqrt{E_{S}}$ is the theoretical sampling errors for the $j$-th photon count bin, which are so small that they are not visible in the output graphics.
IV.3 Numerical scaling
For validation tests to be useful, they must be accurate, efficient and scalable. The results above have validated the accuracy of the positive P-representation by comparing the simulated distribution to an exactly known model. Next, we analyze the efficiency and scalability of simulating GCPs using the positive P-representation for varying mode number $M$ and dimension $d$.
When performing phase-space simulations, one can subdivide the total number of samples into two sub-ensembles $E_{S}=N_{S}N_{R}$. This has both a statistical benefit, as described in the Appendix, and a computational one, as it allows both vector and multi-core parallel computation, drastically reducing computation time.
The first sub-ensemble, $N_{S}$, is the number of samples generated from a phase-space distribution, in this case Eq.(23), corresponding to a specific stochastic trajectory. The number of trajectories is denoted by second sub-ensemble $N_{R}$.
Generally, the size of each sub-ensemble affects the resulting computation time. As $N_{S}\rightarrow\infty$, large memory bandwidth to store vectors and matrices to $64$-bit (double) precision is required. Meanwhile, using $N_{R}\rightarrow\infty$ is more efficient if the hardware has more CPU cores. In practice, so long as $E_{S}$ remains unchanged, one can vary the size of each sub-ensemble to suit hardware requirements without affecting the overall accuracy of the simulation, except that the accuracy of sampling error estimates may be reduced at either extreme.
To investigate the efficiency and scalability of our method, GCPs of dimension $d=1,2,3$ are simulated for GBS networks $M=N=30,60,\dots,480$ with $\boldsymbol{r}=[1,\dots,1]$ and $t=0.6$. Positive-P numerical run-times for these networks are presented in Fig.(2) and were obtained using a typical current (2025) desktop computer with a 12-core 3.7 GHz AMD CPU and 32GB of RAM. All simulations used $E_{S}=2.4\times 10^{6}$ samples, although the sub-ensembles were varied to optimize run-times such that $N_{S}=500,50,5$ and $N_{R}=4.8\times 10^{3},4.8\times 10^{4},4.8\times 10^{5}$ for $d=1,2,3$ respectively. For all networks and dimensions tested, sampling errors were $<10^{-3}$.
From Fig.(2) it is clear that positive-P run-times scale nearly linearly with mode number for all dimensions. As the dimension increases, run-times increase by approximately an order of magnitude. For example, the largest simulated network run-times (in seconds) are $51.34$, $172.3$, and $8800$ for $d=1,2,\>\text{and}\>3$, respectively. This is unexpected, since the floating-point operations needed for the transfer matrix multiplication scale as $M^{2}$, but it indicates that memory access is more important for this range of network sizes.
Figure 2: Run-times in seconds versus mode number of positive-P simulated GCPs of dimension $d=1$ (blue), $d=2$ (orange), and $d=3$ (black). GBS networks with uniform squeezing parameter $\boldsymbol{r}=[1,\dots,1]$ and loss rate $t=0.6$ (i.e. the mean of the parameters used in Fig.(1)) are simulated for mode numbers $M=30,60,\dots,480$ where the network matrix is a Haar random unitary. The algorithm is optimised by only simulating counts within $m_{j}^{(\text{min})}\leq m_{j}\leq m_{j}^{(\text{max})}$. The set of maximum counts for $d=1,2,3$ are $\{62,88,134,218,369\}$, $\{37,50,80,115,190\}$, and $\{25,33,49,77,126\}$ respectively, while the set of minimum counts are $\{0,5,12,45,134\}$, $\{0,0,0,15,60\}$, and $\{0,0,2,11,41\}$. All positive-P simulations are performed using $E_{S}=2.4\times 10^{6}$. Sub-ensemble sizes were varied as $N_{S}=500,50,5$ and $N_{R}=4.8\times 10^{3},4.8\times 10^{4},4.8\times 10^{5}$ for $d=1,2,3$ respectively, in order to optimize run-times on the utilized hardware.
The sharp jump in time from $d=2$ to $d=3$ is due to increased memory requirements. The available memory needed to store a single complex multi-dimensional array to double-precision scales as $16N_{S}(M/d)^{d}$ bytes. Hence when $M=480$, each $d=2$ array is $\approx 0.921N_{S}$ MB, while for $d=3$ it is $\approx 65.5N_{S}$ MB. Naturally, additional memory is necessary to actually perform the computations.
Although more sensitive tests of quantum advantage are possible by comparing $d>1$ GCPs of theory and experiment, there are computational costs. Such costs are not restricted to phase-space representations. Hardware limitations are present in all quantum computing validation schemes. One must therefore balance test sensitivity with available hardware.
@@ V Numerical simulations of experimental networks
In this section, we use the sampled positive-P method to validate experimental data of the GBS device Borealis [5]. The experimental data corresponds to the publicly available dataset of Madsen et al [5]. Other works [46, 47] have validated independently generated data from the Borealis network. These have focused on calculating the quadrature coherence scale [46] and orbits [47] of Borealis.
In this work we focus on quantitatively validating GCPs of different dimensionality, which has not been done previously. Detailed $\chi^{2}$ and Z-statistic tests are applied to compare theory and experiment for both pure and thermalized squeezed state inputs. Due to claims of quantum advantage, most of the analysis in this section is on the count patterns from the $216$-mode high squeezing (HS) parameter and $288$-mode experimental data sets. Summary tables of $\chi^{2}$ and Z-statistic test results for the Borealis data sets can be found in Appendix A.
@@ V.1 Statistical tests
Here we use the shorthand notation $\mathcal{\bar{G}}_{i}=\left\langle\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right\rangle_{P}=E_{S}^{-1}\sum_{e=1}^{E_{S}}\left(\mathcal{G}_{\boldsymbol{S}}^{(n)}(m_{j_{i}})\right)_{(e)}$ for the stochastic trajectory $e\in E_{S}$ to denote the positive-P GCP ensemble means of the $i=1,2,\dots,k$ photon count bin for the $j=1,\dots,d$ grouped count, which converges to the true theoretical GCP in the limit $\mathcal{G}_{S,i}=\lim_{E_{S}\rightarrow\infty}\bar{\mathcal{G}}_{S,i}$ for any general input state $S$. Simulated ensemble means are then compared to the experimental GCPs $\mathcal{G}_{E,i}=N_{E}^{-1}\sum_{i}c_{i}$ formed from $N_{E}\subset\mathcal{S}_{P}$ photon count patterns.
@@
The $\chi^{2}$ tests are used for two reasons: the first is that this is a standard test applied to data from random number generators [16], where the summation is performed over $k$ count bins satisfying $\sum_{i\in S_{j}}c_{i}>10$. The second is that one can include dependence on both theoretical and experimental sampling errors as $\sigma_{i}^{2}=\sigma_{E,i}^{2}+\sigma_{T,i}^{2}$, where $\sigma_{E,i}\propto 1/\sqrt{N_{E}}$ is the estimated experimental sampling error. This provides a method of accurately determining whether observed differences are caused by large sampling error, as is the case for unbinned sparse data.
@@
Output $\chi^{2}$ values follow a distribution that is positively skewed for small $k$. Performing a Wilson-Hilferty transformation [50], which transforms the result as $\chi^{2}\rightarrow\left(\chi^{2}/k\right)^{1/3}$, the resulting distribution converges to a Gaussian distribution $\mathcal{N}\left(\mu_{\mathcal{N}},\sigma_{\mathcal{N}}^{2}\right)$ with mean $\mu_{\mathcal{N}}=(9k-2)/9k$ and variance $\sigma_{\mathcal{N}}^{2}=2/(9k)$ for $k\geq 10$ due to the central limit theorem [50, 51]. This allows one to convert $\chi^{2}$ results into measurable probability distances using approximate Z-statistic, or Z-score, tests [52], where
$Z=\frac{\left(\chi^{2}/k\right)^{1/3}-\mu_{\mathcal{N}}}{\sigma_{\mathcal{N}}}.$ (28)
@@
This measures how many standard deviations away a test statistic is from its expected normally distributed mean. The Z-statistic therefore determines the probability of obtaining the observed result. Agreement within sampling errors produces test results of $\chi^{2}/k\approx 1$ and $\left|Z\right|\leq 1$ when accounting for stochastic fluctuations in the output statistics. For Z-statistic tests we use a regime of validity with the limit $Z\leq 6$ to define accurate comparisons. Above this limit test statistics differ by more than $6\sigma_{\mathcal{N}}$ from their predicted means, corresponding to very small observable probabilities.
The results presented in subsequent subsections use the notation introduced previously [17]. Comparisons between theory and experiment for simulations with pure squeezed states input into the experimental $M\times M$ $\boldsymbol{T}$-matrix are denoted by the subscript $EI$, which we call the ideal ground truth distribution. Our ideal ground truth corresponds to the ground truth distributions used in Ref. [5]. This differs from the non-ideal, or modified, ground truth distribution that includes more realistic experimental effects such as thermalization, and measurement corrections in the theoretical simulations, which are denoted by the subscript $ET$.
The validation tests implemented in this paper focus on comparing experimental data with two theoretical ground truth distributions. If experimental data fail the statistical tests implemented here, it does not necessarily mean there is no possible quantum advantage present in these experiments, but it does mean that the quantum computer has generated the wrong answer relative to the target ground truth distribution.
For example, a result of $Z_{EI}\gg 1$ or $Z_{ET}\gg 1$ for GCPs of dimension $d>1$ indicates high-order correlations present in experimental data are far from their expected values for both tested ground truths. Since quantum advantage requires nonclassical correlations at high-orders, this result may indicate the tested experiment is dominated by low-order correlations, possibly increasing the likelihood of classical sampling algorithms producing samples that are closer to either distribution. The larger the binning dimension, the more sensitive the tests are to these correlations, making them harder to pass. However, at too large a dimension the experimental sampling error increases, reducing the power of the test to discriminate between different distributions.
To determine whether or not there is any quantum advantage requires comparisons with classical sampling algorithms that have a comparable or smaller error. Previous work [17] has performed some of these validation tests for a single sampling algorithm compared to threshold detector experiments. We save a comparable analysis for the PNR experiments considered in this paper for future work.
V.2 $d=1$ grouped count comparisons